In an acute trianGle ABC, an altitude AD is drawn. Find the area of triangle ABC if AB=15 in, AC=18 in, and BD= 10 in.
1
Expert's answer
2020-06-01T18:33:16-0400

Consider the acute triangle as shown in the figure below:





Use Pythagoras theorem in triangle "ABD" as,


"AB^2=BD^2+AD^2"


"15^2=10^2+AD^2"


"225=100+AD^2"


Subtract 100 from both sides to isolate "AD" as,


"225-100=100+AD^2-100"


"125=AD^2"


"AD=\sqrt{125}"


"AD=5\sqrt{5}"


Again, use Pythagoras theorem in triangle "ADC" as,


"AD^2+DC^2=AC^2"


"(5\sqrt{5})^2+DC^2=18^2"


"125+DC^2=324"


Subtract "125" from both sides to isolate "DC" as,


"125+DC^2-125=324-125"


"DC^2=199"


"DC=\sqrt{199}"


So, the length of base of the acute triangle "BC" is,


"BC=BD+DC"


"BC=10+\sqrt{199}"


Now, the area of he triangle is evaluated as,


Area"(A)" "=\frac{1}{2}(BC)(AD)"


"=\frac{1}{2}(10+\sqrt{199})(5\sqrt{5})"


"\approx134.76"


Therefore, the area of the acute triangle is Area"(A)" "=\frac{1}{2}(10+\sqrt{199})(5\sqrt{5})\approx134.76" "in.^2"